# Weak Adversarial Neural Pushforward Method for Fractional Fokker-Planck Equations

### 13 Mar 2026 v1

@@ 5.5 Quantitative Comparison via Robust Statistics
Figure 4: Robust comparison metrics over 50 time snapshots for the fractional Fokker-Planck problem ($\alpha=1.5$): mean (upper left), median (upper right), IQR (center left), MAD (center right), percentile bands (lower left), and standard deviation (lower right). The neural network (solid blue) closely tracks the particle simulation (dashed green) for all robust metrics. The standard deviation panel demonstrates the well-known instability of second-moment statistics for $\alpha<2$ Lévy processes: the particle simulation exhibits large jumps due to occasional extreme samples, while the neural network underestimates the tails. Standard deviation should not be used as a comparison metric for $\alpha<2$.
The mean tracking (upper left panel) shows slightly noisier agreement than the median, as the mean is more sensitive to the tail behavior. Both mean curves nonetheless show the correct monotonic decrease from the initial center $x=1.0$ toward zero, with the discrepancy remaining within normal Monte Carlo sampling variability for K=10, [^0] K=10{,}000 particles.
Taken together, the robust metrics confirm that the learned pushforward distribution matches the particle simulation in location, scale, and spread throughout the time interval $[0,2]$, validating the WANPM approach for fractional Fokker-Planck equations with $\alpha=1.5$.
@@ 6.1 Strengths of the Approach
The extension of WANPM to fractional diffusion is notable for several reasons. First, the computational cost of evaluating the fractional Laplacian on plane-wave test functions is identical to the cost for classical diffusion: a single multiplication by $|w^{(k)}|^{\alpha}$ per sample. This means the method scales to any $\alpha\in(0,2]$ without additional quadrature overhead. Second, the method remains purely mesh-free and requires no spatial discretization, making it in principle applicable to higher-dimensional fFPEs. Third, the pushforward representation naturally handles the heavy-tailed character of the solution by allowing the base distribution to be transformed into a distribution with extended support; the network does not need to represent the density itself, only its moments and cumulative statistics in the weak sense.
@@ 6.2 Limitations and Caveats
Several limitations should be acknowledged. The current experiments are one-dimensional, and scaling to higher dimensions requires further investigation of training stability, the number of test function modes $K$ needed, and the interplay between the base distribution dimension $d$ and the target dimension $n$. In the classical FPE companion paper [3] we have demonstrated scalability to 100 dimensions for standard diffusion; fractional diffusion in high dimensions remains to be explored.
@@
Finally, for $\alpha$ close to 0 the fractional Laplacian becomes nearly singular in frequency space, and the test-function weighting $|w^{(k)}|^{\alpha}$ becomes nearly constant; the adversarial test function optimization may then struggle to identify informative modes. We have not explored this regime and leave it for future work.
@@ 6.3 Relationship to the fBm Fokker-Planck Equation
The implementation notes accompanying this paper also derive a weak formulation for the Fokker-Planck equation driven by fractional Brownian motion (fBm) with Hurst parameter $H\in(0,1)$. In that setting the diffusion operator is non-local in time (due to the memory kernel of fBm) rather than non-local in space. The weak form contains a double integral over the triangular domain $0\leq s\leq t\leq T$, which can be evaluated by a single 3D Monte Carlo without temporal discretization. This is a distinct generalization from the fractional Laplacian case studied here, and we leave its numerical validation to a forthcoming paper.
@@ 6.4 Future Directions
Immediate extensions include: (i) higher-dimensional fFPEs with gradient and non-gradient drifts; (ii) variable-order fractional operators $\alpha(x)$, which can model spatially inhomogeneous anomalous diffusion; (iii) tempered fractional Laplacians, which truncate the heavy tails at a characteristic scale; and (iv) coupling with the fBm memory kernel for subdiffusive systems. On the methodological side, curriculum learning strategies for $\alpha$ (analogous to the $\sigma$-annealing used in our triple-well experiments [3]) may improve training for small $\alpha$.
@@ 7 Conclusion
We have presented a weak adversarial neural pushforward method for the fractional Fokker-Planck equation with fractional Laplacian diffusion of order $\alpha\in(0,2]$. The method combines three key ingredients: (i) a pushforward neural network that enforces the initial condition exactly via a $\sqrt{t}$-prefactored architecture; (ii) plane-wave test functions, which are exact eigenfunctions of the fractional Laplacian so that the operator cost is $\mathcal{O}(1)$ for any $\alpha$; and (iii) a fully Monte Carlo weak-form discretization without temporal mesh, enabling training from samples drawn directly from the product space $0,T]\times\rho_{0}\times\pi_{\text{base}}$.
Numerical validation on a one-dimensional quadratic confining potential with $\alpha=1.5$ demonstrates that the learned distribution accurately tracks the transient evolution over $[0,2]$, as confirmed by robust statistical comparisons with a particle simulation using symmetric $\alpha$-stable Lévy increments. The results also confirm that standard deviation is an unreliable metric for heavy-tailed ($\alpha<2$) distributions and should be replaced by robust alternatives such as IQR and MAD.
The method is mesh-free, requires no density evaluation, no Jacobian computation, and no non-local quadrature, and its computational cost for the fractional operator is identical to that for classical diffusion. These properties make it a promising candidate for solving fractional Fokker-Planck equations in high dimensions.

### 20 Mar 2026 v2

@@ 5.2 Experiment 2: Harmonic Confining Potential
Figure 5: Robust comparison metrics over 50 snapshots, harmonic fFPE ($\alpha=1.5$): mean (upper left), median (upper right), IQR (center left), MAD (center right), percentile bands (lower left), and standard deviation (lower right). Pushforward (solid blue) versus particle simulation (dashed green). The standard deviation panel illustrates the well-known instability of second-moment statistics for $\alpha<2$ and is shown for comparison only; it should not be used as a quantitative metric.
5.3 Experiment 3: Double-Well Potential
Problem setup.
We consider the symmetric double-well potential $V(x)=a(x^{2}-b^{2})^{2}$ with $a=b=1$, which has two minima at $x=\pm 1$ and a barrier of height $V(0)=1$ at $x=0$. The corresponding drift is $b(x)=-V^{\prime}(x)=-4ax(x^{2}-b^{2})=-4x(x^{2}-1)$. The initial condition is a Gaussian centered at the barrier top, $\rho_{0}=\mathcal{N}(0,0.09)$. Under the drift, probability mass flows symmetrically into the two wells while the fractional diffusion generates heavy tails extending well beyond the wells.
This problem is substantially harder than the harmonic case for two reasons. First, the initial condition is located precisely at the local maximum of the potential, so the network must learn a symmetry-breaking dynamics that leads to a bimodal target distribution. A naïve gradient-based method may become trapped in a symmetric unimodal local minimum that satisfies some of the weak-form constraints but fails to capture the double-peak structure. Second, the bimodal steady state has two well-separated mass concentrations that require the pushforward map to simultaneously represent two distinct regions of the target space.
Training configuration.
We use $T=2.0$, K=3, [^0] K=3{,}000 test functions, base dimension $d=8$, and train for 10,000 epochs. The larger $d=8$ (versus $d=5$ for the harmonic case) provides the additional representational capacity needed to model the two-peaked target distribution. All other hyperparameters follow the defaults of Table 1.
Training convergence.
Figure 6 shows the training loss and residual norm over 10,000 epochs. The training here requires substantially more epochs than the harmonic case: the loss exhibits a rapid initial drop but then plateaus at a relatively high value around $10^{0}$$10^{1}$ before gradually descending further. This two-phase behavior is characteristic of bimodal learning: in the first phase, the network learns the approximate location and width of each peak; in the second phase, it refines the peak shapes and the inter-well balance. The adversarial oscillations are more pronounced than in the harmonic case, reflecting the richer structure of the weak-form residuals for a bimodal distribution.
Figure 6: Training history for the double-well fFPE ($\alpha=1.5$, 10,000 epochs). Left: training loss on a log scale. Right: weak-form residual norm on a log scale. Convergence is slower than the harmonic case, reflecting the difficulty of learning bimodal symmetry-breaking dynamics.
Transient distribution comparison.
Figure 7 shows the distribution evolution at 50 uniformly spaced snapshots over $[0,2]$. Starting from the narrow Gaussian at $x=0$, the mass splits symmetrically into the two wells at $x=\pm 1$, progressively forming a bimodal distribution. The pushforward network successfully captures this symmetry-breaking dynamics: the emerging double-peak structure appears at the correct times and with the correct relative peak heights. The progressive accumulation of heavy-tailed mass outside the wells—the fractional diffusion signature—is also visible and reproduced correctly by the network. By $t\approx 0.5$ both peaks are clearly established in the learned distribution, and the subsequent slow drift of the two peaks toward the well minima at $x=\pm 1$ is tracked faithfully until the end of the simulation at $t=2.0$.
Figure 7: Double-well fFPE ($\alpha=1.5$): pushforward samples (blue) versus particle simulation (green) at 50 uniformly spaced time points over $[0,2]$. The potential $V(x)=(x^{2}-1)^{2}$ is shown as a red curve in each panel. The pushforward successfully learns the symmetry-breaking dynamics as mass migrates from the central barrier into both wells simultaneously, together with the development of heavy tails.
5.4 Experiment 4: Triple-Well Potential
Problem setup.
The triple-well potential $V(x)=x^{2}(x^{2}-1)^{2}$ has three local minima at $x\in\{-1,0,1\}$ and two energy barriers at $x\approx\pm 0.577$ of height $V(\pm 1/\sqrt{3})=4/27\approx 0.148$. The drift is $b(x)=-V^{\prime}(x)=-2x(x^{2}-1)(3x^{2}-1)$. The initial condition is an equal-weight bimodal mixture
$\rho_{0}=\tfrac{1}{2}\mathcal{N}(-0.5,0.0225)+\tfrac{1}{2}\mathcal{N}(0.5,0.0225),\quad\sigma_{\mathrm{IC}}=0.15,$ (28)
concentrated near the two inner barriers at $x\approx\pm 0.5$. This initial condition places mass between the inner barriers and the two outer wells, making the subsequent dynamics non-trivial: the network must simultaneously track (i) mass flowing outward over the inner barriers toward the outer wells at $x=\pm 1$, (ii) mass flowing inward toward the central well at $x=0$, and (iii) the development of heavy tails due to fractional diffusion. This is the most challenging of our four benchmark problems.
Training configuration.
We use $T=2.5$, K=2, [^0] K=2{,}000 test functions, base dimension $d=8$, and train for 5,000 epochs. The time horizon is extended to $T=2.5$ to allow the distribution to approach the fractional quasi-steady state, which for the triple-well potential requires longer times than the double-well case due to the three competing potential wells.
Results.
Figure 8 shows the learned distribution (blue) against the particle simulation (orange) at eight time snapshots $t\in\{0.1,0.2,0.5,0.8,1.0,1.5,2.0,2.5\}$. At the earliest time $t=0.1$, the distribution is still primarily bimodal, concentrated near the initial mixture centers at $x\approx\pm 0.5$, and both the learned distribution and the particle simulation agree well at this stage. As time progresses, the dynamics become increasingly complex: at $t=0.2$ small peaks begin to emerge at $x=\pm 1$, and by $t=0.5$ a clear trimodal profile is established, with the central peak at $x=0$ filled in as fractional diffusion drives mass over the inner barriers.
The pushforward network tracks the emergence of all three peaks correctly. By $t=1.0$ the three-peaked profile is well established in both the learned distribution and the particle simulation, with the outer peaks at $x=\pm 1$ dominant and the central peak at $x=0$ smaller but clearly resolved. The correct relative weights of the three peaks and the inter-well heavy-tailed leakage are both reproduced qualitatively throughout. At the longest time $t=2.5$, the distribution has nearly relaxed toward the fractional quasi-steady state, and the agreement between learned and particle distributions remains good across the full support.
Figure 8: Triple-well fFPE ($\alpha=1.5$): WANPM samples (blue) versus particle simulation (orange) at $t\in\{0.1,0.2,0.5,0.8,1.0,1.5,2.0,2.5\}$. The initial bimodal distribution evolves into a trimodal profile as mass redistributes across the three wells, with the central peak at $x=0$ emerging by $t=0.5$ and the distribution approaching the fractional quasi-steady state by $t=2.5$.
5.5 Experiment 5: Steady-State 2D Double-Peak Distribution
Problem setup.
We extend the steady-state framework to two dimensions by considering the potential
$V(x_{1},x_{2})=\bigl[(x_{1}-1)^{2}+(x_{2}-1)^{2}\bigr]\bigl[(x_{1}+1)^{2}+(x_{2}+1)^{2}\bigr],$ (29)
which creates two symmetric potential wells with minima at $(1,1)$ and (1,1)(-1,-1) and a ridge separating them. The drift is $b(x)=-\nabla V(x)$, and the diffusion is isotropic with $\alpha=1.5$. This is the two-dimensional fractional analog of the classical double-well steady-state problem: whereas for $\alpha=2$ the steady-state density is the Boltzmann distribution $\rho_{\infty}\propto e^{-2V(x)/\sigma^{2}}$, for $\alpha=1.5$ no such closed form exists, and the steady-state distribution develops heavy algebraic tails in each direction. The particle benchmark uses $N_{\mathrm{particles}}=10{,}000$ trajectories integrated over $T_{\mathrm{SDE}}=20$ time units.
Training configuration.
The steady-state pushforward network $G_{\bm{\vartheta}}:\mathbb{R}^{8}\to\mathbb{R}^{2}$ maps $\mathbf{r}\sim\mathcal{N}(0,I_{8})$ to samples in $\mathbb{R}^{2}$. We use $K=300$ sinusoidal test functions $f^{(k)}(\mathbf{x})=\sin(\mathbf{w}^{(k)}\cdot\mathbf{x}+b^{(k)})$ with $\mathbf{w}^{(k)}\in\mathbb{R}^{2}$, and a batch size of M=2, [^0] M=2{,}000. Training runs for 10,000 epochs.
Results.
Figure 9 shows scatter plots of 2,000 samples from the learned distribution (left) and from the particle simulation (center), together with the training loss (right). The learned samples cluster around the two potential minima at $(\pm 1,\pm 1)$, correctly reproducing the bimodal structure of the steady-state distribution. Both clusters are elongated and heavy-tailed in agreement with the particle simulation, with the fractional diffusion spreading mass beyond the compact Boltzmann peaks that would arise for $\alpha=2$. The training loss descends from order $10^{1}$ to below $10^{-1}$ within the first 1,000 epochs and continues to decrease slowly, stabilizing around $10^{-1}$ with adversarial oscillations characteristic of the min-max training.
Figure 9: Steady-state fractional 2D double-peak ($\alpha=1.5$): scatter plot of 2,000 learned samples (left, blue) versus particle SDE samples (center, orange), and training loss on a log scale (right). The two clusters at $(\pm 1,\pm 1)$ are correctly identified, with heavy-tailed spread consistent with $\alpha$-stable fractional diffusion.
5.6 Experiment 6: Time-Dependent 2D Ring with Rotational Drift
Problem setup.
We solve the transient fFPE in two dimensions with a ring potential and a divergence-free rotational drift component. The potential is
$V(x_{1},x_{2})=\tfrac{1}{4}(r^{2}-r_{0}^{2})^{2},\qquad r^{2}=x_{1}^{2}+x_{2}^{2},$ (30)
which has a single ring-shaped minimum at radius $r_{0}=2$. The drift combines the radial gradient with a tangential rotation:
$b(x)=-\nabla V(x)+\omega\begin{pmatrix}-x_{2}\\ x_{1}\end{pmatrix}=-2(r^{2}-r_{0}^{2})\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}+\omega\begin{pmatrix}-x_{2}\\ x_{1}\end{pmatrix}$, (31)
with $r_{0}=2$ and angular velocity $\omega=2$. The tangential component $\omega(-x_{2},x_{1})^{T}$ is divergence-free, so it does not affect the steady-state distribution but drives a persistent probability current circulating around the ring. For $\alpha=1.5$, the fractional diffusion spreads mass transversely across the ring with algebraic tails, in contrast to the Gaussian transverse profile that arises for $\alpha=2$. The initial condition is a Gaussian concentrated well inside the ring:
$\rho_{0}=\mathcal{N}\!\left((0,1.2)^{T},\;0.16\,I_{2}\right),$ (32)
and the time interval is $T=0.5$.
Training configuration.
The pushforward network $F_{\bm{\vartheta}}(t,\mathbf{x}_{0},\mathbf{r})=\mathbf{x}_{0}+\sqrt{t}\,\tilde{F}_{\bm{\vartheta}}(t,\mathbf{x}_{0},\mathbf{r})$ maps from $0,T]\times\mathbb{R}^{2}\times\mathbb{R}^{8}$ to $\mathbb{R}^{2}$, using a four-hidden-layer network of width 128 with Tanh activations. We use $K=300$ space-time plane-wave test functions $f^{(k)}(t,\mathbf{x})=\sin(\mathbf{w}^{(k)}\cdot\mathbf{x}+\kappa^{(k)}t+b^{(k)})$ with $\mathbf{w}^{(k)}\in\mathbb{R}^{2}$, batch sizes M=3, [^0] M=3{,}000 (interior) and $M_{0}=M_{T}=1{,}000$ (boundary terms), and train for 5,000 epochs. For the fractional Laplacian acting on these two-dimensional test functions, the eigenfunction property (14) still holds: $(-\Delta_{\mathbf{x}})^{\alpha/2}f^{(k)}=|\mathbf{w}^{(k)}|^{\alpha}f^{(k)}$, where $|\mathbf{w}^{(k)}|$ is the Euclidean norm of the wavenumber vector.
Results.
Figure 10 shows scatter plots of the learned distribution (top row, blue) alongside the particle simulation (bottom row, orange) at six time snapshots $t\in\{0.0$, $0.1$, $0.2$, $0.3$, $0.4$, 0.5}0.5\}. The target ring at radius $r_{0}=2$ is indicated by the red dashed circle. At $t=0$ the initial Gaussian is correctly reproduced, concentrated at $(0,1.2)$ inside the ring. As time evolves, the radial drift carries the mass outward toward the ring while the tangential drift sweeps it counter-clockwise. By $t=0.3$ the samples are already concentrated near the ring, and by $t=0.5$ the ring-supported distribution is well established. The learned distribution closely matches the particle simulation in both the radial concentration at $r\approx r_{0}$ and the angular extent of the distribution at each time. The training loss converges from order $10^{0}$ to below $10^{-2}$ within the first 500 epochs and stabilizes with the adversarial oscillations characteristic of min-max training.
Figure 10: Time-dependent fractional 2D ring with rotational drift ($\alpha=1.5$, $r_{0}=2$, $\omega=2$, $T=0.5$): scatter plots of learned samples (top row, blue) versus particle simulation (bottom row, orange) at $t\in\{0.0,0.1,0.2,0.3,0.4,0.5\}$. The red dashed circle marks the target ring at radius $r_{0}=2$. The learned distribution correctly tracks the outward radial expansion and counter-clockwise rotation throughout.
5.7 Experiment 7: Five-Dimensional Harmonic Potential
Problem setup.
To demonstrate scalability beyond one dimension, we solve the transient fFPE with the isotropic harmonic potential $V(\mathbf{x})=\tfrac{1}{2}\|\mathbf{x}\|^{2}$ in $n=5$ spatial dimensions, with drift $b(\mathbf{x})=-\mathbf{x}$ and $\alpha=1.5$. The initial condition is a tight Gaussian $\rho_{0}=\mathcal{N}(\mathbf{3},0.25\,I_{5})$ (mean $3\mathbf{1}$ in all five dimensions), placed far from the potential minimum at the origin, on the time interval $[0,1]$. This is the five-dimensional fractional analog of Experiment 2: the drift carries each marginal toward zero while the fractional Laplacian develops heavy $\alpha$-stable tails independently in each coordinate. Because the drift and diffusion are isotropic, the marginal distribution in each coordinate evolves identically and should match the one-dimensional fFPE solution.
Training configuration.
The pushforward network maps $(t,\mathbf{x}_{0},\mathbf{r})\in[0,1]\times\mathbb{R}^{5}\times\mathbb{R}^{5}$ to $\mathbb{R}^{5}$ via the $\sqrt{t}$-prefactored architecture (16), with four hidden layers of width 128. We use K=2, [^0] K=2{,}000 plane-wave test functions with wavenectors $\mathbf{w}^{(k)}\in\mathbb{R}^{5}$, batch sizes M=2, [^0] M=2{,}000 (interior) and $M_{0}=M_{T}=1{,}000$ (boundary), and train for 1,500 epochs.
Results.
Figure 11 shows the marginal histograms of the learned distribution (blue) alongside the particle simulation (green) in each of the four displayed dimensions at five time snapshots $t\in\{0.0,0.25,0.5,0.75,1.0\}$. The initial tight Gaussian centered at mean 3 is correctly reproduced at $t=0$ in each dimension. As time evolves, the marginals drift toward zero and develop the characteristic heavy tails of $\alpha$-stable diffusion. The learned marginals closely track the particle simulation histograms throughout, with good agreement in both the bulk location and the tail spread. The behavior is consistent across all five dimensions, confirming that the network has learned a properly symmetric five-dimensional distribution. The training loss decreases from $10^{-1}$ to approximately $3\times 10^{-3}$ and the weak-form residual norm from $10^{1}$ to $2.4$ over 500 epochs, demonstrating rapid and stable convergence. These results confirm that the WANPM framework scales to multi-dimensional fFPEs without modification to the training procedure.
Figure 11: Five-dimensional harmonic fFPE ($\alpha=1.5$, $n=5$, $T=1$, trained for 500 epochs): marginal histograms of the learned distribution (blue) versus particle simulation (green) in dimensions 1–4 at five time snapshots $t\in\{0.0,0.25,0.5,0.75,1.0\}$ (columns). Each row corresponds to one spatial dimension. The initial Gaussian at mean 3 drifts toward the potential minimum at the origin while developing $\alpha$-stable heavy tails consistently across all dimensions; the learned marginals closely track the particle simulation throughout. The omitted bottom panel shows the training loss decreasing from $10^{0}$ to $6.5\times 10^{-3}$ and the weak-form residual norm from $10^{1}$ to $3.69$ over 500 epochs.
@@ 6.1 Strengths of the Approach
The extension of WANPM to fractional diffusion is notable for several reasons. First, the computational cost of evaluating the fractional Laplacian on plane-wave test functions is identical to that for classical diffusion: a single multiplication by $|w^{(k)}|^{\alpha}$ per sample, with no non-local quadrature or dense matrix assembly required. This means the method scales to any $\alpha\in(0,2]$ without additional overhead, and the transition from $\alpha=2$ to $\alpha<2$ requires only a single parameter change in the loss computation.
Second, the method remains purely mesh-free and requires no spatial discretisation, making it in principle applicable to higher-dimensional fFPEs where classical grid methods are prohibitively expensive. Third, the pushforward representation naturally handles the heavy-tailed character of the solution: by representing the distribution implicitly through the pushforward map rather than through an explicit density, the network does not need to resolve the density accurately in the far tail—it only needs to reproduce the distribution correctly in the weak sense, as probed by the plane-wave test functions. This is a significant advantage over PINN or normalizing flow approaches that require pointwise density accuracy.
Fourth, the same framework handles both transient and steady-state problems with minimal modifications. The steady-state problem simply drops the time input and the initial-condition and terminal terms from the loss, as described in Section 2.4. The theoretical result of Proposition 1 provides an exact reference for the fOU steady state, which can serve as a validation benchmark or as the starting point for parameter inference.
6.2 Comparison of the Seven Experiments
The seven experiments exhibit a clear progression in difficulty and dimensionality. The 1D fOU steady state (Experiment 1) is the simplest: no time dependence, a unimodal target with known theoretical form, requiring only 3,000 epochs. The 1D harmonic transient (Experiment 2) adds time dependence but retains a unimodal target; 1,000 epochs suffice. The 1D double-well (Experiment 3) introduces bimodal symmetry-breaking and requires 10,000 epochs and base dimension $d=8$. The 1D triple-well (Experiment 4), though nominally more complex, needs only 5,000 epochs because the bimodal initial condition already distributes mass across the landscape, reducing the symmetry-breaking challenge. The 2D steady-state double-peak (Experiment 5) extends the bimodal problem to two dimensions, requiring 10,000 epochs with $d=8$. The 2D ring (Experiment 6) is the most geometrically complex problem, with a non-gradient rotational drift driving a ring-supported distribution; 5,000 epochs with a four-hidden-layer network prove sufficient. The 5D harmonic (Experiment 7) validates scalability to higher dimensions, where the isotropic structure allows the network to learn a product-form distribution from only 1,500 epochs.
@@ 6.3 Limitations and Caveats
All present experiments are one-dimensional. Scaling to higher dimensions requires further investigation of training stability, the number of test function modes $K$ needed, and the interplay between the base distribution dimension $d$ and the target dimension $n$. In the companion paper [3] we have demonstrated scalability to 100 dimensions for standard diffusion; fractional diffusion in high dimensions remains to be explored. One potential complication is that the $\alpha$-stable tails become increasingly important in high dimensions, and the pushforward map may need a correspondingly larger base dimension $d$ to represent the joint heavy-tailed behavior accurately.
@@
For $\alpha$ close to 0 the fractional Laplacian becomes nearly singular in frequency space and the test-function weighting $|w^{(k)}|^{\alpha}$ becomes nearly constant, potentially making the adversarial optimisation less informative. We have not explored this regime and leave it for future work.
@@ 6.4 Relationship to the fBm Fokker-Planck Equation
The present paper treats anomalous diffusion arising from a fractional Laplacian in space, corresponding to a Lévy-driven SDE. A distinct but related generalisation is the Fokker-Planck equation driven by fractional Brownian motion (fBm) with Hurst parameter $H\in(0,1)$, where the non-locality is in time due to the memory kernel. The weak form in the fBm case contains a double integral over the triangular domain $0\leq s\leq t\leq T$, which can be evaluated by a single 3D Monte Carlo without temporal discretisation. We leave the numerical validation of this distinct generalisation to a forthcoming paper.
@@ 6.5 Future Directions
Immediate extensions include: (i) higher-dimensional fFPEs with gradient and non-gradient drifts, leveraging the scalability demonstrated in [3]; (ii) variable-order fractional operators $\alpha(x)$, which can model spatially inhomogeneous anomalous diffusion; (iii) tempered fractional Laplacians, which truncate the heavy tails at a characteristic scale; and (iv) coupling with the fBm memory kernel for subdiffusive systems. On the methodological side, curriculum learning strategies for $\alpha$—analogous to the $\sigma$-annealing used for the triple-well problem in [3]—may improve training for small $\alpha$ where the test-function weighting becomes uninformative.
@@ 7 Conclusion
We have presented a weak adversarial neural pushforward method for the fractional Fokker-Planck equation with Riesz fractional Laplacian diffusion of order $\alpha\in(0,2]$. The method combines three key ingredients: (i) a pushforward neural network enforcing the initial condition exactly via a $\sqrt{t}$-prefactored architecture (16); (ii) plane-wave test functions that are exact eigenfunctions of the fractional Laplacian, making the fractional operator cost $\mathcal{O}(1)$ for any $\alpha$ (14); and (iii) a fully Monte Carlo weak-form discretisation without temporal mesh, enabling training from samples drawn directly from the product space $0,T]\times\rho_{0}\times\pi_{\mathrm{base}}$.
On the theoretical side, we established the analytical steady-state distribution for the fractional Ornstein–Uhlenbeck process (Proposition 1) by solving the Fourier-space ODE (4), identifying the steady state as the symmetric $\alpha$-stable law $S_{\alpha}((2\theta)^{-1/\alpha},0,\mu)$.
Numerical validation on seven benchmark problems with $\alpha=1.5$ demonstrates that the method accurately captures a wide range of dynamics: relaxation to the $\alpha$-stable fOU steady state; transient evolution in one-dimensional harmonic, double-well, and triple-well potentials; the steady-state bimodal distribution of a 2D double-peak problem; the time-dependent ring-supported distribution of a 2D system with rotational drift; and the isotropic fractional diffusion in five dimensions. Robust statistical comparisons confirm close agreement with particle simulations in all cases, and the results confirm that standard deviation is an unreliable metric for $\alpha<2$.
The method is mesh-free, requires no density evaluation, no Jacobian computation, and no non-local quadrature. Its computational cost for the fractional operator is identical to that for classical diffusion, making it a promising candidate for high-dimensional fractional Fokker-Planck equations.